OpenStudy (anonymous):

does anyone know how to do this? simplify -(4xy)^1/2 (4xy)^3/2

jimthompson5910 (jim_thompson5910):

hint: -a^(1/2)*a^(3/2) = -a^(1/2+3/2) -a^(1/2)*a^(3/2) = -a^(4/2) -a^(1/2)*a^(3/2) = -a^2